Session: Feb 1998 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (tony foley at Wed Feb 25 12:59:40 MET 1998) Minor failures
Initial comments: Formatter clock jump of ~1microsec at 06:50 UT day 56. Telescope steering computer crash ~06:34 to ~07:10 day 56, cause unclear. Observation has occasional EMI, particularly in VC5 LSB.

Onsala (A. Polatidis at Thu Feb 26 09:41:20 MET 1998) Minor failures
The inchword drive of HEAD 2 does not work but HEAD 1 (write head) is working OK Hence the recording data should not be affeceted but there are no readback tests. Lost data between 0119 and 0150 UT (Day 057) due to problems in the last tape change

Westerbork (tony foley at Thu Feb 26 11:16:53 MET 1998) Minor failures
More strange stuff: the tape ran off the end at 07:25 UT on day 56, and the schedule had to be restarted. Clearly at some point the tape was at the wrong place at the wrong time. Schedule was re-started at 07:30 on day 56.

Jodrell (Alastair Gunn at Thu Feb 26 15:03:46 MET 1998) Success
Strong RFI - receiver often saturated. No other known problems.

Torun (K. Borkowski at Thu Feb 26 15:27:10 MET 1998) Success
No known problems

Calibration (jean\-francois Desmurs at Mon Mar 2 18:54:32 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name gc20c.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Thu Mar 5 16:19:20 MET 1998) Failure
A power supply failed shortly after start of experiment at 0655 on first day. This caused formatter to run away. Only first 4 hours of experiment will give useful data.